# A unified harmonic framework for dark siren cosmology

### 13 Mar 2026 v1

@@ Appendix C Cross shot noise
The fundamental assumption of the standard galaxy catalog method is that GW sources are hosted in galaxies. In the cross-correlation method, this is encoded in the cross shot noise: the $1/n^{g,\text{tot}}$ shot noise term in Equation 109, for example, is a direct result from assuming that the GWs are drawn from the galaxies as a doubly stochastic process. Nonetheless, we claimed in the main text that in practice, the cross shot noise does not actually contribute much to the cosmological constraints relative to the information gained from 2-point correlations, due to the binning of the data, a large number of galaxies per bin, and low galaxy completeness. In this section, we make this claim more concrete, as well as show for reference how one could in theory compute the expected cross shot noise for full consistency with the standard galaxy catalog method.
@@
$\langle n^{G_{\alpha}}n^{g_{\beta}}\rangle$ is the Poisson rate of coincidence, which encodes how often a GWs in bin $\alpha$ has a host galaxy in bin $\beta$. For two bins that are perfectly coincident, $\langle n^{G_{\alpha}}n^{g_{\beta}}\rangle=p(D^{g}|z_{\beta})$, the fraction of GW host galaxies observed in the catalog, and thus the shot noise is something like $p(D^{g}|z_{\beta})/n^{g}(z_{\beta})$. This is already quite small if we have a large number of galaxies in each bin or if the galaxy catalog has low completeness, both of which we expect to be true. The cross shot noise is even smaller with arbitrary GW and galaxy bins, which can be computed as
@@
For tophat radial bins, the probability $p^{g}(\beta|z)$ of a galaxy in the catalog with true redshift $z$ to be found in with $\hat{z}\in[\hat{z}_{\min}^{\beta},\hat{z}_{\max}^{\beta}]$ is
@@
$s_{\ell}^{G_{\alpha}g_{\beta}}=\frac{1}{\langle n^{g_{\beta}}\rangle}\int\mathrm{d}zf_{\text{obs}}^{g}(z)\phi^{G_{\alpha}}(z)\int_{\hat{z}_{\min}^{\beta}}^{\hat{z}_{\max}^{\beta}}\mathrm{d}\hat{z}\,\mathcal{L}(\hat{z}|z).$ (127)
@@
The physical explanation of why this is the case is that the identification of a GW host galaxy is difficult, for the standard reasons why the galaxy catalog method is often not informative: galaxy catalogs are large and incomplete, and GW localization errors are large. This is also made worse in the cross-correlation method because we discretize the data and build measurement errors into the theory prediction, such that the GW host requirement can only be imposed up to a certain scale, which is a weaker requirement than the standard galaxy catalog method. Thus, the information on cosmological parameters from matching a GW to a host galaxy is negligible compared to that of the high-SNR measurements of galaxy and GW clustering. It is for these reasons that most multi-tracer applications ignore the cross shot noise altogether, even if the tracers can be coincident (e.g., [109]).
@@ Appendix D BBH population model
In this section, we describe the details of the BBH population model and the implementation of their injection into galaxies . The true values of the population model parameters of can be found in Table 2; recall that we use the Planck 2015 values for $H_{0}$ and $\Omega_{m,0}$ as our fiducial cosmology. We will use $\mathbf{\Lambda}_{c}$, $\mathbf{\Lambda}_{r}$, and $\mathbf{\Lambda}_{m}$ to abbreviate the cosmological, rate, and mass population parameters, respectively.
@@
Thus, we divide the entire underlying set of galaxies into chunks in redshift. For each chunk $i$, we draw the number of galaxies in the chunk $N^{i}$ from a Poisson distribution with expectation $N_{\text{exp}}^{i}$ using Equation 129. Then, we inject $N^{i}$ BBHs into randomly drawn galaxies in the chunk weighted by their redshift as in Equation 128. Finally, for each of the $N_{i}$ BBHs, we draw the primary mass and mass ratio ($m_{1},q$) via rejection sampling, and draw the aligned spins from a uniform distribution $\chi_{1,z},\chi_{2,z}\in[-1,1]$. These parameters, along with with uniformly drawn extrinsic parameters (inclination angle, polarization angle, GPS time of arrival, coalescence phase), are sufficient to pass into gwfast for non-precessing waveforms (including IMRPhenomHM) as mock BBHs. This procedure is repeated for all chunks.
@@
is a Planck taper window to smooth the distribution at minimum mass.
@@
The observed SNRs, posterior measurements, etc. are then computed as described in Section VII.1.3.
@@ Appendix E Quasi-linear corrections to the power spectrum
Figure 9 shows the same corner plot as in Figure 7, but with this added first-order correction parameter $a$; we used a uniform prior of $a\in[-0.2,0.2]$. Compared to the results in Figure 7, we find that adding the nuisance parameter is sufficient to account for unmodeled systematics, removing the bias in $\Omega_{m,0}$ from the galaxy–galaxy analysis. Note that the true values of the nuisance parameters $a=0$ and $b=1.1$ are not recovered exactly correctly, but for our purposes they are nuisance parameters that we marginalize over to encode uncertainty in the shape of the power spectrum. Inferring the shape of the power spectrum requires a more careful analysis that would, for example, include higher-order terms in the Limber approximation.

### 16 Apr 2026 v2

@@ Appendix C Cross shot noise
The fundamental assumption of the standard galaxy catalog method is that GW sources are hosted in galaxies. In the cross-correlation method, this is encoded in the cross shot noise: the $1/n^{g,\text{tot}}$ shot noise term in Equation 113, for example, is a direct result from assuming that the GWs are drawn from the galaxies as a doubly stochastic process. Nonetheless, we claimed in the main text that, in practice, the cross shot noise does not actually contribute much to the cosmological constraints relative to the information gained from 2-point correlations, due to the binning of the data, a large number of galaxies per bin, and low galaxy completeness. In this section, we make this claim more concrete, as well as show for reference how one could in theory compute the expected cross shot noise for full consistency with the standard galaxy catalog method.
@@
$\langle n^{G_{\alpha}}n^{g_{\beta}}\rangle$ is the Poisson rate of coincidence, which encodes how often a GWs in bin $\alpha$ has a host galaxy in bin $\beta$. For two bins that are perfectly coincident, $\langle n^{G_{\alpha}}n^{g_{\beta}}\rangle=p(D^{g}|z_{\beta})\langle n^{G_{\alpha}}\rangle$, the fraction of GW host galaxies observed in the catalog, and thus the shot noise is something like $p(D^{g}|z_{\beta})/n^{g}(z_{\beta})$. This is already quite small if we have a large number of galaxies in each bin or if the galaxy catalog has low completeness, both of which we expect to be true. The cross shot noise is even smaller with arbitrary GW and galaxy bins, which can be computed as
@@
For tophat radial bins, the probability $p^{g}(\beta|z)$ of a galaxy in the catalog with true redshift $z$ to be found in the bin with $\hat{z}\in[\hat{z}_{\min}^{\beta},\hat{z}_{\max}^{\beta}]$ is
@@
$s_{\ell}^{G_{\alpha}g_{\beta}}=\frac{1}{\langle n^{g_{\beta}}\rangle}\int\mathrm{d}z\,p(D^{g}|z)\phi^{G_{\alpha}}(z)\int_{\hat{z}_{\min}^{\beta}}^{\hat{z}_{\max}^{\beta}}\mathrm{d}\hat{z}\,\mathcal{L}(\hat{z}|z).$ (131)
@@
The reason why host galaxies are not so informative are the same as why the standard galaxy catalog method is often not informative: galaxy catalogs are large and incomplete, and secondarily GW localization errors are large. This is also made worse in the cross-correlation method because we discretize the data and build measurement errors into the theory prediction, such that the GW host requirement can only be imposed up to a certain scale, which is a weaker requirement than the standard galaxy catalog method. Thus, the information on cosmological parameters from matching a GW to a host galaxy is negligible compared to that of the high-SNR measurements of galaxy and GW clustering. It is for these reasons that most multi-tracer applications ignore the cross shot noise altogether, even if the tracers can be coincident (e.g., [111]).
@@ Appendix D BBH population model
In this section, we describe the details of the BBH population model and the implementation of their injection into galaxies. The assumed values of the population model parameters of can be found in Table 2; recall that we use the Planck 2015 values for $H_{0}$ and $\Omega_{m,0}$ as our fiducial cosmology. We will use $\mathbf{\Lambda}_{c}$, $\mathbf{\Lambda}_{r}$, and $\mathbf{\Lambda}_{m}$ to abbreviate the cosmological, rate, and mass population parameters, respectively.
@@
We divide the entire underlying set of galaxies into chunks in redshift. For each chunk $i$, we draw the number of BBHs in the chunk $N^{i}$ from a Poisson distribution with expectation $N_{\text{exp}}^{i}$ using Equation 133. Then, we inject $N^{i}$ BBHs into randomly drawn galaxies in the chunk weighted by their redshift as in Equation 132. Finally, for each of the $N_{i}$ BBHs, we draw the primary mass and mass ratio ($m_{1},q$) via rejection sampling, and draw the aligned spins from a uniform distribution $\chi_{1,z},\chi_{2,z}\in[-1,1]$. These parameters, along with uniformly drawn extrinsic parameters (inclination angle, polarization angle, GPS time of arrival, coalescence phase), are sufficient to pass into gwfast for non-precessing waveforms (including IMRPhenomHM) as mock BBHs. This procedure is repeated for all chunks.
@@
is a Planck taper window to smooth the distribution near the minimum mass.
@@
[…]
@@ Appendix E Quasi-linear corrections to the power spectrum
Figure 9 shows the same corner plot as in Figure 7, but with this added first-order correction parameter $a$; we used a uniform prior of $a\in[-0.2,0.2]$. Compared to the results in Figure 7, we find that adding the nuisance parameter is sufficient to account for unmodeled systematics, removing the bias in $\Omega_{m,0}$ from the galaxy–galaxy analysis. Note that the true values of the nuisance parameters $a=0$ and $b=1.1$ are not recovered correctly, but for our purposes they are nuisance parameters that we marginalize over to encode uncertainty in the shape of the power spectrum. Inferring the shape of the power spectrum requires a more careful analysis that would, for example, include higher-order terms in the Limber approximation.